Suppose that Jessica faces the following total cost function: C(y) = 2y2 + y + 200 Note that 200 denotes the fixed cost compoShow that the MC curve intersects the AVC curve and the AC curve at the minimum point.
Answer #1

TC = 2*+ + 2eo MC = ATC 12 АС - чЯ + AC = = гу? +4 H2eo ч AC = 2 + 1 + 2оо AVC = vc - зу? + - зун AFC = - Зоо

DAVC = 2(1) to =a 2 ATC -= 2(1) to - poo = 2-200 DATU ay 2 - 200 - 0 200 y Floo unit y = 10 ATC will be minunn at y =100

AC 75 Arc ° 2 4 6 8 10 12 14 16 18 20 - Quanting As it can be seen ind in the diagram that me curve cuts AC at its minimum po